# Real-Time Recurrent Reinforcement Learning

### v2

@@ Introduction
The last row defines the Jacobian trace recurrently, in terms of the immediate Jacobian and a linear combination of the past trace $\hat{J}_{t}\nabla_{h_{t}}f(x_{t},h_{t})$, allowing to calculate it in parallel to the forward computation of the RNN. When taking an optimization step we can calculate the actual gradients as:
@@
Being online, RTRL is biologically plausible in the time domain. However, the error signal $W^{\top}\varepsilon_{t}$ still assumes weight transport and so do the gradients communicated within the Neural ODE $\hat{J}_{t}\nabla_{h_{t}}f(x_{t},h_{t})$.
In general, RTRL has complexity $\mathcal{O}(n^{4})$ in the number of neurons $n$ compared to $\mathcal{O}(n^{2})$ for BPTT (considering fixed horizon). However, a computationally more efficient version for sparse RNNs was proposed by Menick et al. (2020).
Random-Feedback Local Online Learning (RFLO).
A biologically plausible descendant of RTRL is RFLO (Murray 2019). This leverages the Neural ODE of simple CT-RNNs in order to simplify the RTRL update substantially, dropping all parts that are biologically implausible. RFLO improves biological plausibility of RTRL in two ways:
  1. 1.
    Weight transport during error backpropagation is avoided by using feedback alignment.
  2. 2.
    Locality of gradient information is ensured by dropping all non-local terms from the gradient computation.
@@
where $x_{t}$ is the input at time $t$, and the $1$ concatenated to $\xi_{t}$ accounts for the bias. The output $\hat{y}_{t}\in\mathbb{R}^{O}$ is given by a linear mapping $\hat{y}_{t}=W_{out}h_{t}$. The latent space follows the Neural ODE $\tau\dot{h}=-h_{t}+\varphi(W\xi_{t})$.
RFLO leverages the state-update expression in order to simplify the RTRL update. For brevity, we only show the resulting update rule. A derivation is in the appendix.
Weight transport is avoided by using feedback alignment for propagating gradients. Parameter updates are computed as $\Delta W(t)=\hat{J}_{t}^{W}B\varepsilon_{t}$ using a fixed random matrix $B$. Effective learning is still achievable with this simplified version as shown in (Murray 2019; Marschall, Cho, and Savin 2020).
While RTRL complexity is $\mathcal{O}(n^{4})$ in the number of neurons, RFLO has the same complexity as backpropagation $\mathcal{O}(n^{2})$. The reader is referred to (Murray 2019) and Marschall, Cho, and Savin (2020) for a detailed comparison between RTRL and RFLO.
Algorithm 1 Real-Time Recurrent Reinforcement Learning
0:  Linear policy: $\pi_{\theta_{A}}(a|h)$
0:  Linear value function: $\hat{v}_{\theta_{C}}(h)$
@@
  $\theta_{A},\theta_{C},\theta_{R}\leftarrow$ Randomly initialize parameters
  $B_{A},B_{C}\leftarrow$ Randomly initialize feedback matrices
  $h,e_{A},e_{C},e_{R}\leftarrow\mathbf{0}$
  $o\leftarrow$ Reset Environment
@@
  $v\leftarrow\theta_{C}^{\top}h$
  while not done do
     $\pi\leftarrow\pi_{\theta_{A}}(h)$
     $a\leftarrow$ Sample(dist)
     $o,R\leftarrow$ Take action $a$ in Environment
[…]
     $g_{C}\leftarrow B_{C}\mathbf{1}$triangleright feedback alignment
     $g_{A}\leftarrow B_{A}\nabla_{\text{h}}[\ln\ \pi[a]+\eta_{H}H(\pi)]$
     $e_{C}\leftarrow\gamma\lambda_{C}e_{C}+\alpha h-\alpha\gamma\lambda\left[e_{C}^{\top}h\right]h$
     $e_{A}\leftarrow\gamma\lambda_{A}e_{A}+\nabla_{\theta_{A}}[\ln\ \pi[a]+\eta_{H}H(\pi)]$
     $e_{R}\leftarrow\gamma\lambda_{R}e_{R}+\hat{J}(g_{A}+g_{C})$
     $v^{\prime}\leftarrow\theta_{C}^{\top}h^{\prime}$
     $\delta\leftarrow R+\gamma v^{\prime}-v$triangleright TD error
     $\theta_{C}\leftarrow\theta_{C}+\delta e_{C}+\alpha\left[v-\theta_{C}^{\top}h\right]h$
     $\theta_{A}\leftarrow\theta_{A}+\alpha_{A}\delta e_{A}$
     $\theta_{R}\leftarrow\theta_{R}+\alpha_{R}\delta e_{R}$
     $v\leftarrow v^{\prime},\quad h\leftarrow h^{\prime},\quad\hat{J}\leftarrow\hat{J}^{\prime}$
  end while
@@
Having described each piece of our RTRRL approach, we now proceed discussing how RTRRL puts them together. Algorithm 1 shows the outline of the RTRRL approach. Intuitively, the approximate Jacobian $\hat{J}_{t}$ is computed at each RNN step, and combined with the TD-error of True online TD($\lambda$) to update the RNN weights. We follow the Meta-RL approach delineated in (Wang et al. 2018) feeding in the previous action and reward to the RNN body alongside the observation at each step. $\pi$ and $\hat{v}$ are the Actor and Critic functions parameterized by $\theta_{A}$ and $\theta_{C}$.
@@
We train the Actor and Critic using TD($\lambda$) and take small steps in direction of the action probability and value estimate respectively. The gradients for each function are accumulated by means of additive eligibility traces $e_{A,C}$ with $\lambda$ decay. Additionally, the gradients are passed back to the single layer RNN, parameterized by $\theta_{R}$, through random feedback matrices $B_{A}$ and $B_{C}$ respectively. A third, combined eligibility trace $e_{R}$ is maintained for the RNN, summarizing the combined gradient. For improved exploration, we also compute the gradient of the action distribution’s entropy $H(\pi)$ and add it to the gradient of the policy objective. When using RFLO the passing back of the gradients is done via randomly initialized fixed backwards matrices $B_{A,C}$. This way we arrive at a biologically-plausible algorithm.
@@
The Jacobian in Algorithm 1 can also be computed using RTRL. This reduces the variance in the gradients at a much higher computational cost. Similarly, we can choose to propagate back to the RNN by using the forward weights as in backpropagation. However, we found that the biologically plausible feedback alignment works just as well.
RTRRL Neuron Model.
@@
We use the simplified CT-RNN of the previous section for the RNN body, where we compute the Jacobian $\hat{J}_{t}$ by using RTRL or RFLO as explained above. Extending RFLO, we derive an update rule for the time-constant parameter $\tau$. Again, the full derivation of this Jacobian can be found in Appendix A:
$\hat{J}_{t+1}^{\tau}\approx\hat{J}_{t}^{\tau}(1-\frac{1}{\tau})+\frac{1}{\tau^{2}}\left(h_{t}-\varphi(W\xi_{t})\right)$ (9)
RTRRL Hyperparameters.
The hyperparameters that can be tuned for individual environments are $\gamma$, $\lambda_{A,C,R}$, $\alpha_{A,C}$ and $\alpha$. Our approach does not introduce any new ones over True online TD($\lambda$) other than lambda and learning rate for the RNN. A detailed list of all hyperparameters and RTRRL implementation details are given in the Appendix.
Model Linear TD($\lambda$) RTRL CTRNN RFLO CTRNN PPO CTRNN PPO LSTM
Environment
MemoryChain-4 * 0.07$\pm$0.01 1.00$\pm$0.00 1.00$\pm$0.00 1.00$\pm$0.00 1.00$\pm$0.00
MemoryChain-8 * 0.07$\pm$0.03 1.00$\pm$0.00 1.00$\pm$0.28 1.00$\pm$0.00 1.00$\pm$0.00
MemoryChain-16 * 0.12$\pm$0.04 1.00$\pm$0.36 0.87$\pm$0.42 1.00$\pm$0.00 1.00$\pm$0.00
DeepSea-4 0.99$\pm$0.00 0.99$\pm$0.42 0.99$\pm$0.31 0.99$\pm$0.00 0.99$\pm$0.00
DeepSea-8 0.99$\pm$0.48 0.99$\pm$0.31 0.99$\pm$0.42 0.99$\pm$0.00 0.00$\pm$0.54
DeepSea-16 -0.00$\pm$0.42 0.99$\pm$0.46 0.99$\pm$0.51 0.00$\pm$0.00 0.00$\pm$0.00
Acrobot -500.$0$80.97$\pm$187.54 -86.72$\pm$227.18 -471.40$\pm$221.84 -84.14$\pm$180.65
CartPole-vel * 454.55$\pm$26.52 500.00$\pm$0.00 500.00$\pm$0.00 112.11$\pm$40.38 138.37$\pm$31.41
CartPole-pos * 55.87$\pm$1.18 137.36$\pm$177.76 500.00$\pm$122.30 70.76$\pm$11.42 56.15$\pm$82.74
MetaMaze 15.30$\pm$40.00 17.20$\pm$12.94 19.60$\pm$13.49 10.80$\pm$1.34 10.40$\pm$1.04
DiscountingChain 1.10$\pm$0.05 1.04$\pm$0.04 1.10$\pm$0.04 1.10$\pm$0.00 1.10$\pm$0.00
UmbrellaChain 0.83$\pm$0.31 1.19$\pm$0.59 1.09$\pm$0.60 1.04$\pm$0.12 1.03$\pm$0.12
PointRobot 0.82$\pm$0.51 1.20$\pm$0.42 1.40$\pm$0.27 0.44$\pm$0.05 0.42$\pm$0.07
Reacher 115.47$\pm$17.68 118.70$\pm$8.59 118.77$\pm$10.66 16.72$\pm$0.91 17.27$\pm$1.01
Swimmer 96.56$\pm$4.91 87.69$\pm$10.96 110.72$\pm$62.50 30.76$\pm$1.74 29.05$\pm$1.97
MountCarCont-vel * -41.88$\pm$14.85 -1.31$\pm$13.16 -20.00$\pm$35.18 -1079.70$\pm$167.95 -1126.11$\pm$212.34
MountCarCont-pos * -46.91$\pm$13.71 1.23$\pm$7.27 -59.54$\pm$38.86 -1071.67$\pm$36.35 -1077.49$\pm$114.22
Table 1: Summary of RTRRL experiments: Numbers reported are the median reward and the standard deviation for 5 runs (larger is better), a """""*"" ∗ " marks POMDPs. ”RFLO CTRNN” denotes our biologically plausible instantiation of RTRRL. In each case, the best validation reward achieved throughout training was used. On average, RTRRL arguably achieves the best and most balanced results.
@@
We evaluate the feasibility of our RTRRL approach by testing it on the popular RL benchmarks, provided by the gymnax (Lange 2022) and brax (Freeman et al. 2021) python packages. The tasks comprise fully and partially observable MDPs, with discrete and continuous actions.
The gymnax package implements a range of popular classic and modern control tasks used to benchmark neural-network policies. As baselines we consider TD($\lambda$) with Linear Function Approximation, and Proximal Policy Optimization (PPO) (Schulman et al. 2017), with BPTT for the LSTM and CT-RNN models. Our implementation of PPO is based on purejaxrl (Lu et al. 2022). For each environment, we trained a Meta-RL network with 32 recurrent neurons for either a maximum of 50 million steps or until 20 subsequent epochs showed no improvement.
Our results show that RTRRL with RFLO, as discussed in previous sections, as the default, can often outperform PPO that relies on BPTT. In Table 1 we summarize all our results for the experiments explained in the remainder of this section. The column ”RFLO CT-RNN” corresponds to RTRRL as defined above, but we have also included for comparison, results for RTRRL, where RTRL replaces RFLO. The values presented in the table are the median and standard-deviation of the best evaluation episodes, experienced throughout the training of 5 runs per environment and algorithm.
Discrete Actions.
For discrete actions, the outputs of the actor are the log probabilities for each action, and past actions fed to the RNN are represented in one-hot encoding. In Table 1 we report results for experiments on the classical Acrobot and CartPole benchmarks but also for the bsuite environments (Osband et al. 2020) including DeepSea, MemoryChain and UmbrellaChain, that are specifically tailored to test individual aspects of RL tasks such as exploration and memory capacity. Additionally we included the MetaMaze environments that was introduced by Miconi, Clune, and Stanley (2018) and to test Meta-RL algorithms. As one can see, RTRRL performs best in average, followed by RTRRL with RTRL, as the second best algorithm.
Continuous Actions.
To obtain a stochastic policy, a common trick is to use a probability distribution parameterized by the model output. A natural choice is to use a Gaussian whose mean and standard-deviation are outputs of the neural network. Assuming i.i.d. action components, the network output has to be in $\mathbb{R}^{2U}$ where $U$ is the size of the action vector. An action is obtained by sampling from the distribution. Reparametrization allows for obtaining the probability of the selected continuous action and subsequently computing the gradient of $\log\pi[a]$ in policy-gradient algorithms. In practice, clipping the magnitude of policy gradients to $\leq 1$, in order to avoid exploding gradients that occur when the variance of the action distribution becomes very small, worked best. The continuous action environments included in our experiments, and whose results are illustrated in Table 1 were MountCarCont, PointRobot, Reacher and Swimmer. Moreover, they also included environments implemented in brax, such as, mujoco.
Masked Observations.
We took the classical CartPole MDP and transformed it into a POMDP by masking out parts of the observation. We consider observations that only contain the positions $(x,\theta)$ or the velocities $(\dot{x},\dot{\theta})$, called CartPole-pos and CartPole-vel respectively. The same was done with MountainCarContinuous. Our results show that, when observing positions only, RFLO is the only method able to properly solve this task.
In a similar fashion, we masked observations for the continuous action brax environments. We only kept even entries of the observation, discarding odd ones, to create a POMDP. These physical simulations are computationally demanding which is why we only report the best evaluation epoch for RTRRL after tuning compared to the tuned PPO baselines that were provided by the package authors. The best rewards are $2720.03$ ($2386.09$) for Ant, $2030.10$ ($2317.57$) for Halfcheetah and $1026.78$ ($1072.85$) for Hopper with the tuned baselines in parenthesis, respectively. A plot of the learning curves for the runs in question can be found in Appendix C. As one can see form these results, RTRRL performs on-par with PPO with BPTT, which used batched experiences and handpicked best hyperparameters.
Deep Exploration.
Exploration versus exploitation is a trade-off, central to all agents learning a task online. The DeepSea environment included in bsuite (Osband et al. 2020) is tailored to benchmark the ability of RL algorithms to explore in unfavourable environments. At each step, the agent has to decide on one of two actions which randomly map to left and right in a way that is fixed at each position. A small negative reward is given when the action leading right was chosen and a large positive reward is given at the end of each episode, if the agent chose right each time. The agent is required to explore to reach the rightmost position albeit receiving negative rewards when moving towards it. We give results for exponentially increasing task length.
Figure 4: Boxplots of the Rewards of 10 runs on MemoryChain per type of plasticity for increasing memory lengths. BPTT refers to PPO with LSTM, RFLO and RTRL denote the variants of RTRRL and LocalMSE is a naive approximation to RTRL.
@@
We tested the memory capacity of our RTRRL models by learning to remember the state of a bit for an extended number of steps. The MemoryChain environment can be thought of as a T-maze, known from behavioural experiments with mice. Here, a long corridor leads up to a T-section. Each mouse is given a queue at the entrance to the corridor, indicating on which path leaving the T-section it will find food (Osband et al. 2020).
The experiment tests if the model can remember the state of a bit for a fixed number of time steps. Increasing the number of steps increases difficulty. We conducted MemoryChain experiments with 32 neurons, again with exponentially increasing task length. A boxplot of the results is shown in figure 4. With this experiments we wanted to answer the question of how RTRL, RFLO, and BPTT compare to each-other. The results show that using approximate gradients hampers somewhat memory capacity. Quite surprisingly, RTRRL trained with RTRL, outperformed PPO with LSTMs of the same size that were trained with BPTT.
@@
Ladosz et al. (2022) extended DQNs by adding a neural network head called MOHN, that is subject to Hebbian learning using eligibility traces. Their algorithm significantly outperforms vanilla DQN on confounding POMDPs although slowing down execution time by 66%. However, this work uses back-propagation in its optimization engine.
Ororbia and Mali (2022) introduced a biologically plausible model-based RL framework called active predicitve coding that enables online learning of tasks with extremely sparse feedback. Their algorithm combines predictive coding and active inference, two concepts grounded in neuroscience. Network parameters are trained with Hebbian updates and surrogate losses. This is quite different from our work.
The work of Johard and Ruffaldi (2014) uses a cascade-correlation algorithm with two eligibility traces, that starts from a small neural network and grows it sequentially. This is a very interesting alternative to backpropagation, and it is claimed to be biologically plausible. The two eligibility traces approximate the gradient of the discounted reward, in a connectionist actor-critic setting. The method outperforms regular policy-gradient methods on the CartPole environment. However, it only considers feed-forward networks.
In his Master’s Thesis, Chung (2019) introduces a network architecture similar to RTRRL that consists of a recurrent backbone and linear TD heads. Convergence for the RL algorithm is proven assuming the learning rate of the RNN being magnitudes below those of the heads. Albeit the similarity in network structure, gradients were nonetheless computed using biologically implausible BPTT.
An interesting approach to reduce the complexity of RTRL was proposed by Javed et al. (2023). Similar to the cascade-correlation approach used in Johard and Ruffaldi (2014), they train a recurrent neural network constructively, that is one neuron at a time, keeping preceding neuron’s parameters fixed. This way the RTRL complexity is reduced to the one of BPTT. However, this work does not consider RL.
Recently, Irie, Gopalakrishnan, and Schmidhuber (2023) investigated the practical performance of RTRL-based recurrent RL on a set of memory tasks in 3D. They used LSTMs modified to allow for efficient RTRL updates, and were able to show an improvement over training with BPTT when used in the framework of IMPALA (Espeholt et al. 2018). The main aim of this work was thus to improving RTRL in the context of LSTMs.
A similar idea was explored by Zucchet et al. (2023), who show that independent modules of complex-valued linear recurrent units (LRUs) allow for efficient computation of RTRL updates. Their architecture can be generalized to multi-layer networks setting it apart from other related work. However, their model renounces biological plausibility by restricting the recurrent connections to be diagonal.

### v3

@@ A Meta-RL RNN architecture, resembling the mammalian basal ganglia depicted in figure 1,
Equation 6 defines the Jacobian trace recursively, in terms of the immediate Jacobian and a linear combination of the past trace $\hat{J}_{t}\nabla_{h_{t}}f(x_{t},h_{t})$, allowing to calculate it in parallel to the forward computation of the RNN. When taking an optimization step we can calculate the final gradients as:
@@
Note that RTRL does not require separate phases for forward and backward computation and is therefore biologically plausible in the time domain. However, backpropagating the error signal $W_{out}^{\top}\varepsilon_{t}$ still assumes weight sharing, and the communicated gradients $\hat{J}_{t}\nabla_{h_{t}}f(x_{t},h_{t})$ violate locality.
One big advantage of RTRL-based algorithms, is that the computation time of an update-step is constant in the number of task steps. However, RTRL has complexity $\mathcal{O}(n^{4})$ in the number of neurons $n$ compared to $\mathcal{O}(nT)$ for BPTT, where $T$ is the time horizon of the task. RTRL is therefore not used in practice. Note however, that due to the diagonal connectivity of LRUs, each parameter of their recurrent weight matrix influences just a single neuron. This means that the complexity of the RTRL update is reduced significantly, as shown by Zucchet et al. (2023). Furthermore, they demonstrated how LRUs trained with RTRL can be generalized to the multi-layer network setting. LRU-RNNs are therefore a viable option for efficient online learning.
Random Feedback Local Online Learning (RFLO).
A biologically plausible variant of RTRL is RFLO (Murray 2019). This algorithm leverages the Neural ODE of CT-RNNs in order to simplify the RTRL update substantially, conveniently dropping all parts that are biologically implausible. RFLO improves biological plausibility of RTRL in two ways: 1. Weight transport during error backpropagation is avoided by using FA, 2. Locality of gradient information is ensured by dropping all non-local terms from the gradient computation. Particularly, (7) is leveraged in order to simplify the RTRL update. For brevity, we only show the resulting update rule. A derivation can be found in the Appendix.
$\hat{J}_{t+1}^{W}\approx(1-\tau^{-1})\ \hat{J}_{t}^{W}+\tau^{-1}\varphi^{\prime}(W\xi_{t})^{\top}\xi_{t}$ (8)
Weight transport is avoided by using FA for propagating gradients. Parameter updates are computed as $\Delta W(t)=\hat{J}_{t}^{W}B\varepsilon_{t}$ using a fixed random matrix $B$. Effective learning is still achievable with this simplified version as shown in (Murray 2019; Marschall, Cho, and Savin 2020). RFLO has time complexity $\mathcal{O}(n^{2})$ and is therefore less expensive than BPTT, when the horizon of the task is larger than the number of neurons. The reader is referred to Murray (2019) and Marschall, Cho, and Savin (2020) for a detailed comparison between RTRL and RFLO as well as other approximations.
Algorithm 1 RTRRL
0:  Linear policy: $\pi_{\theta_{A}}(a|h)$
0:  Linear value-function: $\hat{v}_{\theta_{C}}(h)$
@@
1:  $\theta_{A},\theta_{C},\theta_{R}\leftarrow$ initialize parameters
2:  $B_{A},B_{C}\leftarrow$ initialize feedback matrices
3:  $h,e_{A},e_{C},e_{R}\leftarrow\mathbf{0}$
4:  $o\leftarrow$ reset Environment
@@
6:  $v\leftarrow\hat{v}_{\theta_{C}}(h)$
7:  while not done do
8:     $\pi\leftarrow\pi_{\theta_{A}}(h)$
9:     $a\leftarrow$ sample($\pi$)
10:     $o,r\leftarrow$ take action $a$
@@
12:     $e_{C}\leftarrow\gamma\lambda_{C}e_{C}+\nabla_{\theta_{C}}\hat{v}$
13:     $e_{A}\leftarrow\gamma\lambda_{A}e_{A}+\nabla_{\theta_{A}}\log\pi[a]$
14:     $g_{C}\leftarrow B_{C}\mathbf{1}$
15:     $g_{A}\leftarrow B_{A}\nabla_{\pi}\log\ \pi[a]$
16:     $e_{R}\leftarrow\gamma\lambda_{R}e_{R}+\hat{J}(g_{C}+\eta_{A}g_{A})$
17:     $v^{\prime}\leftarrow\hat{v}_{\theta_{C}}(h^{\prime})$
18:     $\delta\leftarrow r+\gamma v^{\prime}-v$
19:     $\theta_{C}\leftarrow\theta_{C}+\alpha_{C}\delta e_{C}$
20:     $\theta_{A}\leftarrow\theta_{A}+\alpha_{A}\delta e_{A}$
21:     $\theta_{R}\leftarrow\theta_{R}+\alpha_{R}\delta e_{R}$
22:     $v\leftarrow v^{\prime},\quad h\leftarrow h^{\prime},\quad\hat{J}\leftarrow\hat{J}^{\prime}$
23:  end while
@@
Figure 3: Boxplot of the combined normalized validation rewards achieved for 5 runs each on a range of different environments from the gymnax, and popgym packages. Depicted are results for RTRRL with CTRNNs and LRUs, each with and without FA. RTRRL-LRU-Meta and PPO-CTRNN-Meta perform best overall. Using FA always leads to diminished performance. Fully biologically plausible RTRRL-RFLO with FA often achieves on-par results.
@@
Having described all the necessary pieces, we now discuss how RTRRL puts them together. Algorithm 1 shows the outline of the RTRRL approach. Importantly, on line 11, the next latent vector $h^{\prime}$ is computed by the single-layer RNN parameterized by $\theta_{R}$. We use the Meta-RL structure depicted in figure 1. Therefore, previous action, reward, and observation are concatenated to serve as input $[o,a,r]$. The approximate Jacobian $\hat{J}_{t}$ is computed as second output of the RNN step, and is later combined with the TD-error $\delta$ and eligibility trace $e_{R}$ of TD($\lambda$) to update the RNN weights. $\pi$ and $\hat{v}$ are the Actor and Critic functions parameterized by $\theta_{A}$ and $\theta_{C}$. We train the Actor and Critic using TD($\lambda$) and take small steps in direction of the log of the action probability $\pi[a]=\mathbf{P}\mathcal{N}(\theta_{A}^{\top}h^{\prime})$ and value estimate $\hat{v}=\theta_{V}^{\top}h^{\prime}$ respectively. The gradients for each function are accumulated using eligibility traces $e_{A,C}$ with $\lambda$ decay. Additionally, the gradients are passed back to the RNN through random feedback matrices $B_{A}$ and $B_{C}$ respectively. The eligibility trace $e_{R}$ for the RNN summarizes the combined gradient. The use of randomly initialized fixed backwards matrices $B_{A,C}$ in RFLO makes RTRRL biologically plausible.
@@
The Jacobian in Algorithm 1 can also be computed using RTRL. This reduces the variance in the gradients at a much higher computational cost. Similarly, we can choose to propagate back to the RNN by using the forward weights as in backpropagation. However, we found that the biologically plausible Feedback Alignment works in many cases.
@@
The fully connected shared RNN layer acts as a filter in classical control theory, and estimates the latent state of the POMDP. Linear Actor and Critic functions then take the latent state as input and compute action and value respectively. We use CT-RNNs or LRUs as introduced in the previous section for the RNN body, where we compute the Jacobian $\hat{J}_{t}$ by using RFLO or RTRL as explained above. Extending RFLO, we derive an update rule for the time-constant parameter $\tau$. The full derivation can be found in the Appendix:
$\hat{J}_{t+1}^{\tau}\approx\hat{J}_{t}^{\tau}(1-\tau^{-1})+\tau^{-2}\left(h_{t}-\varphi(W\xi_{t})\right)$ (9)
The hyperparameters of RTRRL are $\gamma$, $\lambda_{[A,C,R]}$, $\alpha_{[A,C,R]}$. Our approach does not introduce any new ones over TD($\lambda$) other than lambda and learning rate for the RNN. For improved exploration, we also compute the gradient of the action distribution’s entropy, scaled by a factor $\eta_{H}$, and add it to the gradients of policy and RNN. In order to stay as concise as possible, this is omitted from algorithm 1. Further implementation details can be found in the Appendix.
Figure 4: Left: Boxplots of 10 runs on MemoryChain per type of plasticity for increasing memory lengths. BPTT refers to PPO with LSTM, RFLO and RTRL denote the variants of RTRRL and LocalMSE is a naive approximation to RTRL. Right: Tuning the entropy rate is a trade-off of best possible reward vs. consistency as shown for the MetaMaze environment.
Figure 5: Left: Shown are the mean rewards aggregated over 5 runs each; shaded regions are the variance. While in many cases it does not make a difference, not using the Meta-RL architecture hampers performance in some cases. Using biologically plausible Feedback Alignment can lead to worse results, but more often than not it does not have a significant impact.
@@
We evaluate the feasibility of our RTRRL approach by testing on RL benchmarks provided by the gymnax (Lange 2022), popgym  (Morad et al. 2022) and brax (Freeman et al. 2021) packages. The tasks comprise fully and partially observable MDPs, with discrete and continuous actions. As baselines we consider TD($\lambda$) with Linear Function Approximation, and Proximal Policy Optimization (PPO) (Schulman et al. 2017) using the same RNN models but trained with BPTT. Our implementation of PPO is based on purejaxrl (Lu et al. 2022). We used a truncation horizon of 32 for BPTT. For each environment, we trained a network with 32 neurons for either a maximum of 50 million steps or until 20 subsequent epochs showed no improvement. The same set of hyperparameters, given in the Appendix, was used for all the RTRRL experiments if not stated otherwise. Importantly, a batch size of $1$ was used to ensure biological plausibility. All $\lambda$’s and $\gamma$ were kept at $0.99$, $\eta_{H}$ was set to $10^{-5}$, and the adam (Kingma and Ba 2015) optimizer with a learning rate of $10^{-3}$ was used.
For discrete actions, the outputs of the actor are the log probabilities for each action, and past actions fed to the RNN are represented in one-hot encoding. To obtain a stochastic policy for continuous actions, a common trick is to use a Gaussian distribution parameterized by the model output. We can then easily compute the gradient of $\log\pi[a]$.
Figure 3 shows a selection of our experimental results as boxplots. Depicted are the best validation rewards of 5 runs each. RTRRL-LRU achieves the best median reward in almost all cases, outperforming PPO. Interestingly, Linear TD($\lambda$) was able to perform very well on some environments that did require delayed credit-assignment, such as UmbrellaChain. At the same time it performed especially poor on POMDP environments such as StatelessCartPole. Finally, fully biologically plausible RTRRL-RFLO often shows comparable performance to the other methods and too can outperform PPO in some cases. One notable exception is NoisyStatelessCartPole, where RTRRL-RFLO performed best. This could hint at RFLO being advantageous in noisy environments. Investigating this theory however was outside of the scope of this paper. A second exception is RepeatPrevious, where RTRRL-RFLO performed surprisingly poorly for unknown reasons. Finally, we include results for both RTRRL versions when used with FA.
@@
We compared the memory capacity of RTRRL-CTRNN by learning to remember the state of a bit for an extended number of steps. The MemoryChain environment can be thought of as a T-maze, known from behavioral experiments with mice. (Osband et al. 2020) The experiment tests if the model can remember the state of a bit for a fixed number of time steps. Increasing the number of steps increases the difficulty. We conducted MemoryChain experiments with 32 neurons, for exponentially increasing task lengths. A boxplot of the results is shown in figure 4. With this experiments we wanted to answer the question of how RTRL, RFLO, and BPTT compare to each-other. The results show that using approximate gradients hampers somewhat memory capacity. Quite surprisingly, RTRRL outperformed networks of the same size that were trained with BPTT.
Figure 6: Rewards over training epochs for environments from the brax package, masked to make them POMDP. Shown are the best run for RTRRL vs. a tuned PPO baseline. RTRRL performance shows increased variance due to the batch size of 1.
Ablation Experiments.
In order to investigate the importance of each integral part of RTRRL, we conducted three different ablation studies with CT-RNNs. Note that for the first two ablation experiments, $\eta_{H}$ was kept at 00.
  • Feedback Alignment: We conducted experiments to test whether using the biologically plausible FA would result in a decreased reward. While true for most environments, surprisingly, FA often performs just as well as regular backpropagation, as can be seen in figures 3 and 5.
  • Meta-RL: We tried out RTRRL without the Meta-RL architecture and found that some tasks, such as RepeatPrevious, gain a lot from using the Meta-RL approach, as depicted in figure 5. However, the results show that for many environments it does not matter.
  • Entropy rate: One tunable hyperparameter of RTRRL is the weight given to the entropy loss $\eta_{H}$. We conducted experiments with varying magnitudes and found, that the entropy rate represents a trade-off between consistency and best possible reward. In other words, the entropy rate seems to adjust the variance of the resulting reward, as depicted in figure 4 (right).
Physics Simulations.
We masked the continuous action brax environments, keeping even entries of the observation and discarding odd ones, to create a POMDP. For each environment, we ran hyperparameter tuning for at least 10 hours and picked the best performing run. Figure 6 shows the evaluation rewards for RTRRL, compared to the tuned PPO baselines that were provided by the package authors.
@@
Johard and Ruffaldi (2014) used a cascade-correlation algorithm with two eligibility traces, that starts from a small neural network and grows it sequentially. This interesting approach, featuring a connectionist actor-critic architecture, was claimed to be biologically plausible, although they only considered feed-forward networks and the experimenta evaluation was limited to the CartPole environment.
In his Master’s Thesis, Chung (2019) introduced a network architecture similar to RTRRL that consists of a recurrent backbone and linear TD heads. Convergence for the RL algorithm is proven assuming the learning rate of the RNN is magnitudes below those of the heads. Albeit using a network structure similar to RTRRL, gradients were nonetheless computed using biologically implausible BPTT.
Ororbia and Mali (2022) introduced a biologically plausible model-based RL framework called active predicitve coding that enables online learning of tasks with extremely sparse feedback. Their algorithm combines predictive coding and active inference, two concepts grounded in neuroscience. Network parameters were trained with Hebbian updates combined with component-wise surrogate losses.
One approach to reduce the complexity of RTRL was proposed by Javed et al. (2023). Similar to Johard and Ruffaldi (2014), a RNN is trained constructively, one neuron at a time, subsequently reducing RTRL complexity to the one of BPTT. However, this work did not consider RL.
Recently, Irie, Gopalakrishnan, and Schmidhuber (2024) investigated a range of RNN architectures for which the RTRL updates can be computed efficiently. They assessed the practical performance of RTRL-based recurrent RL on a set of memory tasks, using modified LSTMs, and were able to show an improvement over training with BPTT when used in the framework of IMPALA (Espeholt et al. 2018). Noteworthy, IMPALA requires collecting complete episodes before computing updates making it biologically implausible.